# Quantum resonances and scattering poles of classical rank one locally symmetric spaces

### v1

@@ 5. Relating scattering matrices and resolvent kernels
The key to the comparison between scattering matrices and resolvent kernels for global symmetric spaces given in [HHP19] was the possibility to take boundary values of Laplace eigenfunctions on $X$. The boundary value maps are given as the inverses of Poisson transforms. Moreover, the proof of [BO12, Lemma 11] suggests a way how to use the knowledge about the relation between resolvent poles and scattering poles of rank one global symmetric spaces in order to determine the corresponding relation for rank one locally symmetric spaces. We will follow this strategy and start by reviewing a few facts about Poisson transforms and boundary value maps.
@@ 5.1. Poisson transforms and boundary values
$\beta_{\rho-\lambda}P_{\lambda}=c(\lambda)\mathrm{Id}.$
@@
$\beta_{1,\lambda}\circ P^{\gamma}_{\lambda}=c_{\gamma}(\lambda)\,\mathrm{Id}.$
The definitions of the $c$-functions in [HHP19] and [Olb94] are compatible. In our situation $\sigma$ and $\gamma$ are trivial, which means that the boundary value map $\beta_{1,\lambda}$ from [Olb94] equals the boundary value map $\beta_{\rho-\lambda}$ from [HHP19].
@@
In principle the construction of the boundary value maps is built on the study of asymptotic behavior of eigenfunctions. The construction in [KO77] depends on a specific and very technical theory, the algebraic analysis initiated by Sato. If the boundary value map were applied to moderately growing eigenfunctions on all of $X\times X$ we could use [vdBS87] to describe the boundary value maps. This description shows that the direct product situation leads to a decoupling of the asymptotics which leads to a decoupling of the boundary value map. Taking into account the notation and the various conventions from Remarks 5.1 and 4.3, the boundary value map from [BO12, § 6.3] suitably restricted is given by
@@
$\beta_{\nu}=\beta_{\nu(\zeta)}:=\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}\otimes\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}},\qquad\zeta\in\mathcal{H}_{\mathrm{phys}},$ (5.2)
@@
with the boundary value map of [HHP19] on the right-hand side. and $\nu\in(\mathfrak{a}_{\mathbb{C}}^{*})^{2}=\mathbb{C}^{2}$ denoting the leading exponents of the eigensections one takes boundary values of.
We restrict our attention to the non-vanishing boundary value maps. For these the parameter $\nu\in\mathbb{C}^{2}$ is given by
$\nu:=\nu(z):=(-\left\|\rho\right\|-\lambda(z),-\left\|\rho\right\|-\lambda(z)),$
where $\lambda(z):=\sqrt{\left\|\rho\right\|^{2}-z}\in\{\mathrm{Re}\,>0\}\subset\mathbb{C}$ (c.f. Remark 4.3), $z\in\mathbb{C}$ is a Casimir eigenvalue, and we use the convention from [BO12, § 6.1] that the complex vector space ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ is identified with $\mathbb{C}$ by complex linear extension of the unique order-preserving isometric isomorphism ${\bf\mathfrak{a}}^{\ast}\cong\mathbb{R}$. This means that $\rho\in{\bf\mathfrak{a}}^{\ast}$ is identified with the number $\left\|\rho\right\|\in\mathbb{R}$ (which accordingly is written simply as $\rho$ in [BO12, § 6]), where the norm is with respect to the inner product on ${\bf\mathfrak{a}}^{\ast}\cong{\bf\mathfrak{a}}$ given by restriction of the chosen bilinear form $b$ on ${\bf\mathfrak{g}}$ defining the Riemannian metric on $X=G/K$. In particular, the number $\lambda(z)\in\mathbb{C}$ corresponds to the element
$\lambda(z)=\sqrt{\left\|\rho\right\|^{2}-z}\,\frac{\rho}{\left\|\rho\right\|}=\sqrt{1-\frac{z}{\left\|\rho\right\|^{2}}}\,\rho\in{\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}.$ (5.3)
The number $-\left\|\rho\right\|-\lambda(z)$ is a leading exponent of a Casimir eigenfunction for the eigenvalue $z$. The boundary value map maps tensor products of Casimir eigenfunctions for the eigenvalue $z$ to tensor products of elements of $H^{1,\lambda(z)}$ (see [BO12, Lemmas 10 & 11] for all this and take the change of signs described in Remark 4.2 into account). Thus $\beta_{\nu}$ coincides with $\beta_{1,-\lambda(z)}\otimes\beta_{1,-\lambda(z)}$, where $\beta_{1,\lambda}$ is as in [Olb94], so that in view of (ii) we have
$\beta_{\nu}=\beta_{\rho+\lambda(z)}\otimes\beta_{\rho+\lambda(z)}$ (5.4)
with the boundary value maps from [HHP19] on the right hand side and the parameters considered as elements of ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$.
To express the relation (5.4) using parameters in $\mathbb{C}$, one needs to take into account that in [HHP19] the identification of ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ with $\mathbb{C}$ differs from that in [BO12, § 6]: In [HHP19] the element $\rho\in{\bf\mathfrak{a}}^{\ast}$ gets identified with the number $\frac{\left\|\rho\right\|}{\left\|\alpha_{0}\right\|}\in\mathbb{R}$, where $\alpha_{0}\in\Sigma^{+}\subset{\bf\mathfrak{a}}^{\ast}$ is the unique reduced positive root and the norm is with respect to the Killing form. Let us choose the bilinear form $b$ in [BO12, § 6.1] such that on ${\bf\mathfrak{a}}\times{\bf\mathfrak{a}}$ it agrees with the Killing form, so that the norms on ${\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ considered in [BO12, § 6.1] and [HHP19] agree. Furthermore, let us denote for any element $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ the corresponding complex numbers in [BO12, § 6] and [HHP19] by $\lambda_{\mathrm{BO}}\in\mathbb{C}$ and $\lambda_{\mathrm{HHP}}\in\mathbb{C}$, respectively. Then
$\lambda_{\mathrm{BO}}=\left\|\alpha_{0}\right\|\lambda_{\mathrm{HHP}}.$ (5.5)
In particular, for the element $\lambda(z)\in{\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ from (5.3), which satisfies $\lambda(z)_{\mathrm{BO}}=\sqrt{\left\|\rho\right\|^{2}-z}=\sqrt{\rho_{\mathrm{BO}}^{2}-z}$, we find
$\begin{split}\lambda(z)_{\mathrm{HHP}}&=\frac{1}{\left\|\alpha_{0}\right\|}\sqrt{\left\|\rho\right\|^{2}-z}=\sqrt{\bigg{(}\frac{\left\|\rho\right\|}{\left\|\alpha_{0}\right\|}\bigg{)}^{2}-\frac{z}{\left\|\alpha_{0}\right\|^{2}}}=\sqrt{\rho_{\mathrm{HHP}}^{2}-\frac{z}{\left\|\alpha_{0}\right\|^{2}}}.\end{split}$ (5.6)
@@
Let $f,g\in{C^{\infty}}(\partial X)$ have disjoint supports. Then applying [HHP19, last eq. on p. 20] with the parameter ${-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ gives
$\left\langle s_{\zeta},f\otimes g\right\rangle$ $=\left\langle S_{\zeta}f,g\right\rangle$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}(\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}R_{\zeta})^{\prime}f,g\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}R_{\zeta}\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime}f,\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime}g\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}r_{\zeta}|_{X\times X\setminus\mathrm{diag}_{X}},(\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime}\otimes\beta_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}^{\prime})(f\otimes g)\big{>}$
@@ 5.2. Local-global comparisons
Recall the Schwartz kernels $r_{\zeta}$ and $\tilde{r}^{\Gamma}_{\zeta}$ of $R_{\zeta}$ and $\tilde{R}^{\Gamma}_{\zeta}$, respectively from Section 3. Further, let $s_{\zeta}$ and $\tilde{s}^{\Gamma}_{\zeta}$ be the Schwartz kernels of $S_{\zeta}$ and the $\Gamma$-invariant lifts of the Schwartz kernel of $S^{\Gamma}_{i\zeta\left\|\alpha_{0}\right\|}$, respectively. We consider all of these as meromorphically extended to $\mathbb{C}$.
@@
$\beta_{\nu(\zeta)}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}})=a_{\zeta}\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}.$
Actually, [BO12, p. 140] does only yield Equation (5.9) on a neighborhood of $\Omega\times\Omega\setminus(D_{\Gamma}^{\Omega}\setminus\mathrm{diag}_{\Omega})$ in the closure of $X\times X$ (intersected with $X\times X\setminus\mathrm{diag}_{X})$. But this suffices as the boundary values only depend on the values on a neighborhood of the boundary.
@@
Outside the set $-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$, the scattering poles which are not in the physical plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ are precisely the quantum resonances.
@@
Let $\zeta_{0}$ be a quantum resonance of $\Gamma\backslash X$, i.e. a pole of $\tilde{r}_{\zeta}^{\Gamma}$. Then $\zeta_{0}$ is not in the physical halfplane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ and Lemma 5.3 implies that $\zeta_{0}$ is also a pole of $\tilde{r}_{\zeta}^{\Gamma}|_{X\times X\backslash D_{\Gamma}}$.
Case 1 $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$ and $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}\not=0$:
@@
Case 2 $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$ and $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}=0$:
@@
In view of [HHP19, Re. 5.1] $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}=0$ implies $\mathrm{Im}\,\zeta_{0}>\|\rho\|$. If $\delta_{\Gamma}<2\|\rho\|$, then $i\rho\in\mathcal{H}^{\Gamma}_{\mathrm{phys}}$, hence also $\zeta_{0}\in\mathcal{H}^{\Gamma}_{\mathrm{phys}}$. Thus, this case cannot occur, i.e. $\delta_{\Gamma}=2\|\rho\|$. But then 00 is in the $L^{2}$-spectrum of $\Delta_{\Gamma\backslash X}$. Since the continuous $L^{2}$-spectrum of $\Delta_{\Gamma\backslash X}$ is $[\|\rho\|^{2},\infty[$ the point 00 has to be an eigenvalue. But $\Delta_{\Gamma\backslash X}f=0$ for $f\in L^{2}(\Gamma\backslash X)$ is equivalent to $df=0$ and $d^{*}f=0$, so $f$ has to be constant. But this implies that $\Gamma\backslash X$ has finite volume, which contradicts our hypothesis the $\Gamma$ is convex-cocompact, but not cocompact.
@@
Conversely, assume that $\zeta_{0}$ is a scattering pole which is not in the physical halfplane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ and satisfies $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$. Then, by the argument given in Case 2, $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}\not=0$ and again Equation (5.2) and [HHP19, Thm. 6.4] imply that $\zeta\mapsto\beta_{\nu(\zeta)}$ is a holomorphic family of injective operators in a neighborhood of $\zeta_{0}$.
@@
For any $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ and $f\in C^{-\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ [BO12, p. 134] yields an Eisenstein series, which we denote by $\mathcal{P}_{\lambda}^{\Gamma}f$ since it takes the role of the Poisson transform in the case of locally symmetric spaces. Then $\mathcal{P}_{\lambda}^{\Gamma}f\in C^{\infty}(\Gamma\backslash X)$ is a $\Delta_{\Gamma\backslash X}$-eigenfunction, which has $f$ as boundary value. Applying this to the product $\Gamma\backslash\Omega\times\Gamma\backslash\Omega$ and scattering kernel $\tilde{s}_{\zeta}^{\Gamma}$ with the appropriate $\lambda$ determined by $\zeta$ we find a kernel function $r^{\Gamma,\mathrm{Eis}}_{\zeta}$ on $\Gamma\backslash X\times\Gamma\backslash X$ which has the same boundary values as $r^{\Gamma}_{\zeta}$. But then, since $\zeta\mapsto\beta_{\nu(\zeta)}$ in a neighborhood of $\zeta_{0}$ is a holomorphic family of injective maps, we have $r^{\Gamma,\mathrm{Eis}}_{\zeta}=r^{\Gamma}_{\zeta}$ and $\zeta_{0}$ is a pole of $r^{\Gamma}_{\zeta}$, i.e. it is also a quantum resonance. ∎

### v2

@@ 5. Relating scattering matrices and resolvent kernels
The key to the comparison between scattering matrices and resolvent kernels for global symmetric spaces given in [HHP19] was the possibility to take boundary values of Laplace eigenfunctions on $X$. The boundary value maps are given by the inverses of Poisson transforms. Moreover, the proof of [BO12, Lemma 11] suggests a way how to use the knowledge about the relation between resolvent poles and scattering poles of rank one global symmetric spaces in order to determine the corresponding relation for rank one locally symmetric spaces. We will follow this strategy and start by reviewing a few facts about Poisson transforms and boundary value maps.
@@ 5.1. Poisson transforms and boundary values
$\beta^{\mathrm{HHP}}_{\rho-\lambda}\circ P_{\lambda}=c(\lambda)\mathrm{Id},$
where we decorated the boundary value map with HHP in order to avoid confusion later on.
@@
$\beta^{\mathrm{Olb94}}_{1,\lambda}\circ P^{\gamma}_{\lambda}=c_{\gamma}(\lambda)\,\mathrm{Id}.$
The definitions of the $c$-functions in [HHP19] and [Olb94] are compatible. In our situation $\sigma$ and $\gamma$ are trivial, which means that the boundary value map $\beta^{\mathrm{Olb94}}_{1,\lambda}$ from [Olb94] equals the boundary value map $\beta^{\mathrm{HHP}}_{\rho-\lambda}$ from [HHP19].
@@
In principle the construction of the boundary value maps is built on the study of asymptotic behavior of eigenfunctions. The construction in [KO77] depends on a specific and very technical theory, the algebraic analysis initiated by Sato. When the boundary value map is applied to moderately growing eigenfunctions on all of $X\times X$ we can use [vdBS87] to describe the boundary value maps. This description shows that the direct product situation leads to a decoupling of the asymptotics which leads to a decoupling of the boundary value map. Taking into account the notation and the various conventions from Remarks 5.1 and 4.4, and using the first of the isomorphisms (4.8), the boundary value map from [BO12, § 6.3] suitably restricted is given by
@@
$\zeta\in\mathcal{H}_{\mathrm{phys}}:\qquad\beta_{\nu}=\beta_{\nu(\zeta)}:=\beta^{\mathrm{HHP}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}\otimes\beta^{\mathrm{HHP}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}},$ (5.2)
@@
with $\nu\in(\mathfrak{a}_{\mathbb{C}}^{*})^{2}\cong\mathbb{C}^{2}$ denoting the leading exponents of the eigensections one takes boundary values of.
Recall the Schwartz kernels $r_{\zeta}$ and $\tilde{r}^{\Gamma}_{\zeta}$ of $R_{\zeta}$ and $\tilde{R}^{\Gamma}_{\zeta}$, respectively from Section 3. Further, let $s_{\zeta}$ and $\tilde{s}^{\Gamma}_{\zeta}$ be the Schwartz kernels of $\mathbf{S}_{\zeta}$ and the $\Gamma$-invariant lift of the Schwartz kernel of $\mathbf{S}^{\Gamma}_{\zeta}$, respectively. We consider all of these as meromorphically extended to $\mathbb{C}$.
@@
Let $f,g\in{C^{\infty}}(\partial X)$ have disjoint supports. Then recalling (5.2) and applying [HHP19, last eq. on p. 20] with the parameter ${-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ gives
$\left\langle s_{\zeta},f\otimes g\right\rangle$ $=\left\langle\mathbf{S}_{\zeta}f,g\right\rangle$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}\beta^{\mathrm{HHP}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}(\beta^{\mathrm{HHP}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}R_{\zeta})^{\prime}f,g\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}R_{\zeta}\beta^{\mathrm{HHP}^{\prime}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}f,\beta^{\mathrm{HHP}^{\prime}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}g\big{>}$
$=-2i\left\|\alpha_{0}\right\|\zeta\big{<}r_{\zeta}|_{X\times X\setminus\mathrm{diag}_{X}},(\beta^{\mathrm{HHP}^{\prime}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}\otimes\beta^{\mathrm{HHP}^{\prime}}_{\rho-i\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}})(f\otimes g)\big{>}$
@@ 5.2. Local-global comparisons
$\beta_{\nu(\zeta)}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}})=a_{\zeta}\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}.$ (5.6)
Actually, [BO12, p. 140] does only yield Equation (5.5) on a neighborhood of $\Omega\times\Omega\setminus(D_{\Gamma}^{\Omega}\setminus\mathrm{diag}_{\Omega})$ in the compactification of $X\times X$ (intersected with $X\times X\setminus\mathrm{diag}_{X})$. But this suffices as the boundary values only depend on the values on a neighborhood of the boundary.
@@
Outside the set $-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}_{0}$, the scattering poles which are not in the physical plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ are precisely the quantum resonances.
Moreover, if $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}_{0}$ is a simple pole of the quantum resolvent as well as the scattering matrix, then the boundary value map induces an isomorphism between the images of the two residue operators at $\zeta_{0}$.
@@
Let $\zeta_{0}$ be a quantum resonance of $\Gamma\backslash X$, i.e. a pole of $\tilde{r}_{\zeta}^{\Gamma}$. Then $\zeta_{0}$ is not in the physical halfplane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ and Lemma 5.3 implies that $\zeta_{0}$ is also a pole of $\tilde{r}_{\zeta}^{\Gamma}|_{X\times X\backslash D_{\Gamma}}$. Now suppose that $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}_{0}$.
Case 1 $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}\not=0$:
@@
Case 2 $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}=0$:
@@
In view of [HHP19, Re. 5.1] $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}=0$ implies $\mathrm{Im}\,\zeta_{0}>\|\rho\|$. If $\delta_{\Gamma}<2\|\rho\|$, then $i\rho\in\mathcal{H}^{\Gamma}_{\mathrm{phys}}$, hence also $\zeta_{0}\in\mathcal{H}^{\Gamma}_{\mathrm{phys}}$. Thus, this case cannot occur, i.e. $\delta_{\Gamma}=2\|\rho\|$. But then 00 is in the $L^{2}$-spectrum of $\Delta_{\Gamma\backslash X}$. Since the continuous $L^{2}$-spectrum of $\Delta_{\Gamma\backslash X}$ is $[\|\rho\|^{2},\infty[$ the point 00 has to be an eigenvalue. But $\Delta_{\Gamma\backslash X}f=0$ for $f\in L^{2}(\Gamma\backslash X)$ is equivalent to $df=0$ and $d^{*}f=0$, so $f$ has to be constant. But this implies that $\Gamma\backslash X$ has finite volume, which contradicts our hypothesis that $\Gamma$ is convex-cocompact, but not cocompact.
@@
Conversely, assume that $\zeta_{0}$ is a scattering pole which is not in the physical halfplane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ and satisfies $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}_{0}$. Then, by the argument given in Case 2, $c\big{(}i\zeta_{0}\frac{\alpha_{0}}{\left\|\alpha_{0}\right\|}\big{)}\not=0$ and again Equation (5.2) and [HHP19, Thm. 6.4] imply that $\zeta\mapsto\beta_{\nu(\zeta)}$ is a holomorphic family of injective operators in a neighborhood of $\zeta_{0}$.
@@
For any $\lambda\in{\bf\mathfrak{a}}_{\mathbb{C}}^{\ast}$ and $f\in C^{-\infty}(\Gamma\backslash\Omega,V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ [BO12, p. 134] yields an Eisenstein series, which we denote by $\mathcal{P}_{\lambda}^{\Gamma}f$ since it takes the role of the Poisson transform in the case of locally symmetric spaces. Then $\mathcal{P}_{\lambda}^{\Gamma}f\in C^{\infty}(\Gamma\backslash X)$ is a $\Delta_{\Gamma\backslash X}$-eigenfunction, which has $f$ as boundary value. Applying this to the product $\Gamma\backslash\Omega\times\Gamma\backslash\Omega$ and scattering kernel $\tilde{s}_{\zeta}^{\Gamma}$ with the appropriate $\lambda$ determined by $\zeta$ we find a kernel function $r^{\Gamma,\mathrm{Eis}}_{\zeta}$ on $\Gamma\backslash X\times\Gamma\backslash X$ which has the same boundary values as $r^{\Gamma}_{\zeta}$. But then, since $\zeta\mapsto\beta_{\nu(\zeta)}$ in a neighborhood of $\zeta_{0}$ is a holomorphic family of injective maps, we have $r^{\Gamma,\mathrm{Eis}}_{\zeta}=r^{\Gamma}_{\zeta}$ and $\zeta_{0}$ is a pole of $r^{\Gamma}_{\zeta}$, i.e. it is also a quantum resonance.
To prove the claim about the resonance multiplicities, let $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}_{0}$ be a quantum resonance. Then we are in Case 1 above, so that $\zeta\mapsto\beta_{\nu(\zeta)}$ is a holomorphic family of injective operators in a neighborhood of $\zeta_{0}$. If $\zeta_{0}$ is a simple pole of $\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}}$, then the residue of $\beta_{\nu(\zeta)}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}})$ at $\zeta_{0}$ is given by
$\mathrm{Res}_{\zeta=\zeta_{0}}\big{(}\beta_{\nu(\zeta)}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}})\big{)}=\beta_{\nu(\zeta_{0})}\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}}).$
On the other hand, we have
$\mathrm{Res}_{\zeta=\zeta_{0}}(a_{\zeta}\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}})=\frac{i}{2\left\|\alpha_{0}\right\|}\mathrm{Res}_{\zeta=\zeta_{0}}\big{(}\frac{1}{\zeta}\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}\big{)}.$
If $\zeta_{0}$ is also a simple pole of $\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}$, then
$\frac{i}{2\left\|\alpha_{0}\right\|}\mathrm{Res}_{\zeta=\zeta_{0}}\big{(}\frac{1}{\zeta}\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}\big{)}=\frac{i}{2\left\|\alpha_{0}\right\|\zeta_{0}}\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}).$
Together, this gives
$\beta_{\nu(\zeta_{0})}\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{r}^{\Gamma}_{\zeta}|_{X\times X\setminus D_{\Gamma}})=\frac{i}{2\left\|\alpha_{0}\right\|\zeta_{0}}\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{s}_{\zeta}^{\Gamma}|_{\Omega\times\Omega\setminus D^{\Omega}_{\Gamma}}).$ (5.18)
We want to use this relation between kernel residues to get a relation between the images of the operator residues. To this end, we first observe that for all $\varphi,\psi\in{C^{\infty}_{c}}(X)$
$\langle\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{r}^{\Gamma}_{\zeta}),\varphi\otimes\psi\rangle$ $=\mathrm{Res}_{\zeta=\zeta_{0}}(\langle\tilde{r}^{\Gamma}_{\zeta},\varphi\otimes\psi\rangle)$
$=\mathrm{Res}_{\zeta=\zeta_{0}}(\langle\tilde{R}^{\Gamma}_{\zeta}\varphi,\psi\rangle)$
$=\langle\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{R}^{\Gamma}_{\zeta})\varphi,\psi\rangle.$
Thus, $\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{r}^{\Gamma}_{\zeta})$ is the Schwartz kernel of $\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{R}^{\Gamma}_{\zeta})$. Similarly, $\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{s}^{\Gamma}_{\zeta})$ is the Schwartz kernel of $\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{\mathbf{S}}^{\Gamma}_{\zeta})$, where $\tilde{\mathbf{S}}^{\Gamma}_{\zeta}:C^{\infty}_{c}(\partial X)\to\mathcal{D}^{\prime}(\partial X)$ is the linear operator with Schwartz kernel $\tilde{s}^{\Gamma}_{\zeta}$. To continue, let us use the simplified notation $\tilde{\mathsf{R}}:=\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{R}^{\Gamma}_{\zeta})$, $\tilde{\mathsf{S}}:=\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{\mathbf{S}}^{\Gamma}_{\zeta})$, $\tilde{\mathsf{r}}:=\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{r}^{\Gamma}_{\zeta})$, $\tilde{\mathsf{s}}:=\mathrm{Res}_{\zeta=\zeta_{0}}(\tilde{s}^{\Gamma}_{\zeta})$ as well as the somewhat abstract notation $\mathrm{dom}(\mathsf{X})$ for the domain of an operator $\mathsf{X}$. Then we have
$\mathrm{im}(\tilde{\mathsf{R}})=\mathrm{span}\{\tilde{\mathsf{r}}(f\otimes\cdot)\,|\,f\in\mathrm{dom}(\tilde{\mathsf{R}})\},$ (5.19)
while
$\mathrm{im}(\tilde{\mathsf{S}})$ $=\mathrm{span}\{\tilde{\mathsf{s}}(\omega\otimes\cdot)\,|\,\omega\in\mathrm{dom}(\tilde{\mathsf{S}})\}$
=(5.18)${{=}}\mathrm{span}\{\beta_{\nu(\zeta_{0})}(\tilde{\mathsf{r}})(\omega\otimes\cdot)\,|\,\omega\in\mathrm{dom}(\tilde{\mathsf{S}})\}$ ) }
=(5.2)${{=}}\mathrm{span}\{\beta^{\mathrm{HHP}}_{\rho-i\frac{\zeta_{0}\alpha_{0}}{\left\|\alpha_{0}\right\|}}\big{(}\tilde{\mathsf{r}}(\beta^{\mathrm{HHP}^{\prime}}_{\rho-i\frac{\zeta_{0}\alpha_{0}}{\left\|\alpha_{0}\right\|}}(\omega)\otimes\cdot)\big{)}\,|\,\omega\in\mathrm{dom}(\tilde{\mathsf{S}})\}$ ) }
$=\mathrm{span}\{\beta^{\mathrm{HHP}}_{\rho-i\frac{\zeta_{0}\alpha_{0}}{\left\|\alpha_{0}\right\|}}\big{(}\tilde{\mathsf{r}}(f\otimes\cdot)\big{)}\,|\,f\in\mathrm{dom}(\tilde{\mathsf{R}})\}$
=(5.19)${{=}}\beta^{\mathrm{HHP}}_{\rho-i\frac{\zeta_{0}\alpha_{0}}{\left\|\alpha_{0}\right\|}}\big{(}\mathrm{im}(\tilde{\mathsf{R}})\big{)},$
where in the second last step we used that the adjoint $\beta^{\mathrm{HHP^{\prime}}}_{\rho-i\frac{\zeta_{0}\alpha_{0}}{\left\|\alpha_{0}\right\|}}:\mathrm{dom}(\tilde{\mathsf{S}})\to\mathrm{dom}(\tilde{\mathsf{R}})$ is surjective (being the adjoint of an injective map). Now, analogously as at the end of the proof of [HHP19, Thm. 7.1], we observe that as a map $\mathrm{im}(\tilde{\mathsf{R}})\to\mathrm{im}(\tilde{\mathsf{S}})$ the boundary value map is inverted by the Poisson transform not only from the right but also from the left, so that it is bijective. Thus, passing to $\Gamma$-quotients using the $\Gamma$-equivariance of the boundary value map, we arrive at a linear isomorphism
$\mathrm{im}(\mathrm{Res}_{\zeta=\zeta_{0}}(\mathbf{S}_{\zeta}^{\Gamma}))\cong\mathrm{im}(\mathrm{Res}_{\zeta=\zeta_{0}}(R_{\zeta}^{\Gamma})).$
This finishes the proof. ∎